8. Solve the differential equation y+y=sin(2) cos(x).
Answer #1

We can solve the given initial value problem by using method of variation of parameters method.

So the most general solution is   Y = YC + YP

YC = Homogeneous solutin

YP = Particular solution.

Now the given initil value problem is y’’ + y = sinx cosx

So let us take the homogeneous equation for solving

So y’’ + y = 0    -------------------(4)

Now let us find the characteristic equation for homogeneous equation by assuming the solution y = erx not equla to Zero.

So y’ = r erx

And y’’ = r2erx

so substitute in (1)

y’’ + y = 0

r2erx + erx = 0

erx (r2+1) = 0

erx is not equal to zero so r2+1 = 0

so r2+1 = 0

r2  = -1

r2  = i2

r1 = +i and r2 = -i  =>  0 +/- i   in the form of a +/- ib so a = 0 and b = 1

so when the characteristic equation roots are complex and imaginary then homogeneous solution is in the form as given below.

Yc = C1eax cos(bx)+ C2eax sin(bx)

Yc = C1e(0)x cos(x)+ C2e(0)x sin(x)

Yc = C1 cos(x)+ C2 sin(x)

So now we need to find the wronskian for particular solution

Yc = C1y1(x) + C2y2(x)

So y1(x) = cosx and y2(x) = sinx

y1’(x) = - sinx and y2’(x) = cosx

And  W(y1 , y2) = y1(x) y2’(x) - y1’(x) y2(x)

W(y1 , y2) = (cosx) (cosx) – (-sinx) (sinx)

W(y1 , y2) = (cos2x) + (sin2x)

W(y1 , y2) = 1

And non homogeneous term g(x) = sinx cosx

solution So the particular up a nyong f 4,13).911) dx Y2(al. gral dat Yraw w14, 42) Baroul wly, 92) suin. Seen. Csa Con j upa